OpenStudy (anonymous):

The sum of the two numbers is 12. the difference of the same two numbers is 4. What is the larger of the two numbers?

OpenStudy (anonymous):

x + y = 12 x-y = 4 solve for x on one, then use substitution

OpenStudy (p0sitr0n):

yes, y=12-x x-(12-x)=4 -12+2x=4 2x=16 x=8 y=4 so 8